OpenStudy (anonymous):

Englebert wants to buy a new car, and has narrowed his choices down to 2 possibilities. The first possibility is a hybrid that gets 100 miles per gallon gas mileage, but costs $35,000. The second possibility is a sedan that gets 25 miles a gallon but only costs $20,000. Englebert is going to pay cash for either car. He drives 25,000 miles a year, in an area where the gas price is $4.00 a gallon. •For both the hybrid and sedan cases, write linear equations that show what Englebert's cost of ownership is versus the years he owns the cars. Work collectively with your classmates on this pr

OpenStudy (anonymous):

do you know how to solve systems of linear equations?

OpenStudy (anonymous):

Not that much.......still learning......can you give example.....would help.....not looking for just answer....but help...

OpenStudy (anonymous):

OK. Well let's take the first car. It gets 100 miles per gallon. He drives 25,000 miles a year. So how many gallons does he use in a year?

OpenStudy (anonymous):

2500000

OpenStudy (anonymous):

That would be right if the car needed 100 gallons of gas to go one mile. But it is the opposite: it goes 100 miles with 1 gallon of gas.

OpenStudy (anonymous):

ooohhhh ok 1000

OpenStudy (anonymous):

how do you get 1000?

OpenStudy (anonymous):

1000 is not correct

OpenStudy (anonymous):

help

OpenStudy (anonymous):

If you can go 100 miles with one gallon, how many miles can you go with two gallons?

OpenStudy (anonymous):

200

OpenStudy (anonymous):

And how many miles with 3 gallons?

OpenStudy (anonymous):

300

OpenStudy (anonymous):

OK ... so if you want to go 500 miles, how many gallons do you need?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

and if you want to go 5,000 miles, how many gallons?

OpenStudy (anonymous):

1000

OpenStudy (anonymous):

why 1000?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the relationship is that if we have X number of gallons, we can go 100x miles.

OpenStudy (anonymous):

50

OpenStudy (anonymous):

And similarly, if we want to go Y miles, we need Y/100 gallons. right?

OpenStudy (anonymous):

So how many gallons if we want to go 25,000 miles?

OpenStudy (anonymous):

(50 was right, for 5,000 miles!)

OpenStudy (anonymous):

250

OpenStudy (anonymous):

perfect!!! OK, so we know in 1 year he uses 250 gallons of gas. And gas costs $4 per gallon. so how much money does he spend on gas in a year?

OpenStudy (anonymous):

1000

OpenStudy (anonymous):

yes! And that's in one year, so if he drives the car for, say, "Y" years, how much does he spend on gas?

OpenStudy (anonymous):

what is Y

OpenStudy (anonymous):

the number of years he drives the car

OpenStudy (anonymous):

So your answer can have the "Y" in it

OpenStudy (anonymous):

ok sorry for being so difficult....

OpenStudy (anonymous):

So it would be C=

OpenStudy (anonymous):

the cost of the car plus the yearly price of gas, both added together

OpenStudy (anonymous):

C=gc + miles C=1000+35,000 C=45,000

OpenStudy (anonymous):

well gas isn't 1000 total, it's 1000 per year, so 45000 is only correct if he were driving for one year. But the question asked you for an equation of the total cost versus number of years. If he's driving for Y years, then the cost of the hbyrid is C= 35,000 + 1000Y Now, you can do the same thing for the other car, too

OpenStudy (anonymous):

•How many years will it take for the gas savings with the hybrid to make up for the larger purchase price of the hybrid?

OpenStudy (anonymous):

First you have to find the equation for the other car.

OpenStudy (anonymous):

ok so the other car would be C= one moment

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

I know to divide 25,000 miles by 25 miles a gallon.....then what?

OpenStudy (anonymous):

that tells you how many gallons he uses. multiply that by cost per gallon.

OpenStudy (anonymous):

that's all per year, remember

OpenStudy (anonymous):

40000

OpenStudy (anonymous):

4000......sorry

OpenStudy (anonymous):

per year, right? so what's the equation for the total cost as a function of Y, the number of years?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what's the equation for the total cost as a function of Y, the number of years?

OpenStudy (anonymous):

C=25,000 +4000Y or C=25,000+1000Y

OpenStudy (anonymous):

very very close - remember the two cars cost different amounts

OpenStudy (anonymous):

you got the hard part which is the cost of gass per year, but you didn't get the initial price of the car

OpenStudy (anonymous):

C=20,000 +4000Y

OpenStudy (anonymous):

right, and what about the other one?

OpenStudy (anonymous):

C=35,000 +1000Y

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

•How many years will it take for the gas savings with the hybrid to make up for the larger purchase price of the hybrid

OpenStudy (anonymous):

they're asking how many years does it take before the total cost of the two cars are equal. so set the two equations equal to each other and solve for Y.

OpenStudy (anonymous):

? Help plzzzz...you are so helpful

OpenStudy (anonymous):

35,000 +1000Y = 20,000 +4000Y

OpenStudy (anonymous):

So I need to get the Y's isolated

OpenStudy (anonymous):

exactamundo

OpenStudy (anonymous):

right

OpenStudy (anonymous):

35,000+1000Y=20,000+4000Y -40000Y -4000Y 35,000+3000Y=20,000Y -

OpenStudy (anonymous):

So far am I right?

OpenStudy (anonymous):

You are on the right track but got the sign wrong on the left side. 1000 minus 4000 isn't +3000.

OpenStudy (anonymous):

It's probably easier to move the 1000Y instead of moving the 4000Y (but, the answer will come out the same no matter which way you choose)

OpenStudy (anonymous):

30,000Y=20,000............so lost right now....

OpenStudy (anonymous):

35,000+1000Y=20,000+4000Y Subtract 1000Y from both sides. Then subtract 20,000 from both sides. What do you get?

OpenStudy (anonymous):

35,000Y=23,000Y

OpenStudy (anonymous):

no, let's do it step by step. First let's subtract 1000Y. 35,000 + 1000Y = 20,000 + 4000Y - 1000Y - 1000Y What do you get?

OpenStudy (anonymous):

35,000y =23,000Y I have problems with my negative and positive

OpenStudy (anonymous):

You're mixing up the terms that have Y and the terms that do not have Y. If you're subtracting a Y term you ignore the terms that don't have Ys.

OpenStudy (anonymous):

100 + 35Y - 35Y is just equal to 100.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So what's 35,000 + 1000Y - 1000Y?